# A vectorial Darboux transformation for integrable matrix versions of the Fokas-Lenells equation

### v1

@@ 6.2.3 The double root case
Let $n=1$. If (6.50) admits a single root $r_{1}$ and a double root $r_{2}$, then it is equivalent to
@@
Since $\alpha_{1}\neq\alpha_{2}$, the imaginary parts vanish if and only if $\mathrm{Re}(r_{2})=0$. Hence
$r_{2}=\mathrm{i}\,\kappa\,,\qquad\kappa\in\mathbb{R}\,$,
@@
$|A_{1}|^{2}$ $=$ $\frac{(\alpha_{2}-2\kappa)\,(2k\kappa-1-2\alpha_{1}k)^{2}}{4\alpha_{1}^{2}\,k\,(\alpha_{2}-\alpha_{1})}\,,\qquad|A_{2}|^{2}=-\frac{(\alpha_{1}-2\kappa)\,(2k\kappa-1-2\alpha_{2}k)^{2}}{4\alpha_{2}^{2}\,k\,(\alpha_{2}-\alpha_{1})}\,$,
@@
$\frac{\alpha_{2}-2\kappa}{k\,(\alpha_{2}-\alpha_{1})}>0\,,\qquad\frac{\alpha_{1}-2\kappa}{k\,(\alpha_{2}-\alpha_{1})}<0\,$,
and implies $(\alpha_{1}-2\kappa)(\alpha_{2}-2\kappa)<0$. The Lyapunov equation can only be satisfied if either $c_{1}=0$ or $c_{2}=c_{3}=0$.
@@
$(A^{\dagger}\boldsymbol{\alpha}A)^{2}\,\Big{(}|c_{3}|^{2}+\frac{\|A\|^{4}\,|\tilde{c}|^{2}}{|A_{1}|^{2}\,|A_{2}|^{2}\,(\alpha_{2}-\alpha_{1})^{2}}\Big{)}=k\,\|A\|^{2}\,\Big{|}\mathrm{i}\,c_{2}-c_{3}\,\kappa+c_{3}\,\big{(}\frac{1}{2k}+\frac{A^{\dagger}\boldsymbol{\alpha}A}{\|A\|^{2}}\big{)}-\tilde{c}\,\Big{|}^{2}\,$,
@@
$\tilde{c}$ $=$ $\Big{(}(1-A^{\dagger}\boldsymbol{\alpha}A)\,k^{-1}+\frac{\alpha_{1}\alpha_{2}\|A\|^{2}}{A^{\dagger}\boldsymbol{\alpha}A}\Big{)}^{-1}\Big{(}2\mathrm{i}\,c_{2}\,\kappa-c_{3}\,\kappa^{2}-\mathrm{i}\,\frac{A^{\dagger}\boldsymbol{\alpha}^{2}A}{A^{\dagger}\boldsymbol{\alpha}A}(c_{2}+\mathrm{i}\,c_{3}\,\kappa)+\frac{1}{4}k^{-2}c_{3}$
@@
$\Omega$ $=$ $\frac{\|A\|^{2}}{k(A^{\dagger}\boldsymbol{\alpha}A)^{2}\,\varrho_{2}^{2}}\Big{(}\frac{\mathrm{i}\,k\,\tilde{r}\,|c_{2}|^{2}}{3\alpha_{1}\alpha_{2}}\varrho_{3}\,(\mathrm{i}\,A^{\dagger}\boldsymbol{\alpha}A\,\varrho_{2}-\varrho_{3})\,(3xt^{2}-\frac{3\mathrm{i}\,\alpha_{1}\alpha_{2}}{k\tilde{r}}x^{2}t+\frac{\mathrm{i}\,k\tilde{r}}{\alpha_{1}\alpha_{2}}t^{3})$
$-\mathrm{i}\,\big{(}k|c_{2}|^{2}A^{\dagger}\boldsymbol{\alpha}A\,\varrho_{2}(\varrho_{2}-k\varrho_{1})+2\mathrm{i}\,\mathrm{Re}(c_{2}c_{3}^{\ast})\,\varrho_{3}(\mathrm{i}\,A^{\dagger}\boldsymbol{\alpha}A\,\varrho_{2}-\varrho_{3})\big{)}\,(xt+\frac{\mathrm{i}\,k\tilde{r}}{2\alpha_{1}\alpha_{2}}t^{2})$
@@
$+\mathrm{i}\,k|c_{2}|^{2}A^{\dagger}\boldsymbol{\alpha}A\,\|A\|^{2}\varrho_{4}^{2}\,(k\varrho_{1}-\varrho_{2})\,\big{(}\mathrm{i}\,k\kappa\,(k\varrho_{1}-\varrho_{2})+\varrho_{3}\big{)}+\varrho_{8}\big{)}\,x+c\Big{)}\,$,
@@
$\tilde{r}$ $=$ $\mathrm{i}\,(\alpha_{1}+\alpha_{2}-2\kappa)\,$,
$\varrho_{1}$ $=$ $-2\,\kappa\,A^{\dagger}\boldsymbol{\alpha}A+A^{\dagger}\boldsymbol{\alpha}^{2}A\,$,
@@
$\varrho_{3}$ $=$ $-\mathrm{i}\,k^{2}\,\big{(}A^{\dagger}\boldsymbol{\alpha}A\,\kappa^{2}-(\alpha_{1}\alpha_{2}\,\|A\|^{2}+A^{\dagger}\boldsymbol{\alpha}^{2}A)\,\kappa+\alpha_{1}\alpha_{2}\,A^{\dagger}\boldsymbol{\alpha}A\big{)}$
$-\mathrm{i}\,k\,\big{(}\kappa\,(A^{\dagger}\boldsymbol{\alpha}A)^{2}+\frac{1}{2}\alpha_{1}\alpha_{2}\|A\|^{2}+\frac{1}{2}\varrho_{1}\big{)}+\frac{1}{4}\mathrm{i}\,A^{\dagger}\boldsymbol{\alpha}A\,(2A^{\dagger}\boldsymbol{\alpha}A-1)\,$,
$\varrho_{4}$ $=$ $\frac{1}{2k^{2}}\Big{(}\|A\|^{2}\,\big{(}2\mathrm{i}\,(\mathrm{i}\,k\kappa\,\varrho_{2}+\varrho_{3})+\varrho_{2}\big{)}+2kA^{\dagger}\boldsymbol{\alpha}A\,\varrho_{2}\Big{)}\,$,
$\varrho_{5}$ $=$ $(A^{\dagger}\boldsymbol{\alpha}A)^{2}\varrho_{2}\,(1-2k\kappa)\,(k\varrho_{4}-A^{\dagger}\boldsymbol{\alpha}A\,\varrho_{2})-2\alpha_{1}\alpha_{2}\|A\|^{2}(k(A^{\dagger}\boldsymbol{\alpha}A)^{2}\varrho_{2}^{2}+\|A\|^{2}\varrho_{3}^{2})\,$,
$\varrho_{6}$ $=$ $\|A\|^{2}\varrho_{1}\,(1-2k\kappa)-2(k\varrho_{4}-A^{\dagger}\boldsymbol{\alpha}A\,\varrho_{2})\,$,
$\varrho_{7}$ $=$ $(\alpha_{1}\alpha_{2}\|A\|^{2}-\kappa\,A^{\dagger}\boldsymbol{\alpha}A)\,(k\,(A^{\dagger}\boldsymbol{\alpha}A)^{2}\varrho_{2}^{2}+\|A\|^{2}\varrho_{3}^{2})\,$,
$\varrho_{8}$ $=$ $|c_{2}|^{2}A^{\dagger}\boldsymbol{\alpha}A\,\|A\|^{4}\varrho_{1}\varrho_{3}^{2}\varrho_{4}+\frac{1}{2}\varrho_{4}\,\Big{(}4\mathrm{i}\,\mathrm{Im}(c_{2}c_{3}^{\ast})\,\big{(}k\,\kappa\,A^{\dagger}\boldsymbol{\alpha}A\,\|A\|^{2}\varrho_{3}\varrho_{4}\,(k\varrho_{1}-\varrho_{2})+\mathrm{i}\,\|A\|^{2}\varrho_{1}\varrho_{7}\big{)}$
$+\mathrm{i}\,k\,c_{2}^{\ast}c_{3}\,(A^{\dagger}\boldsymbol{\alpha}A)^{2}\|A\|^{2}\varrho_{1}\varrho_{2}\varrho_{4}\,(1-2k\kappa)-|c_{3}|^{2}\varrho_{4}\varrho_{5}\Big{)}\,$,
@@
Figure 12: Plots of the absolute values of the two components of a positon-type solution from the class in Example 6.9 (2). Here we set $\alpha_{1}=1,\alpha_{2}=2,k=1,\kappa=3/4,c_{1}=0,c_{2}=1+\mathrm{i}\,$ and $c_{3}=1+(353/1365)\,\mathrm{i}\,$, so that the constraint from the Lyapunov equation is satisfied.
@@
Next we present a “second-order” version of the above class of solutions in Example 6.9.
@@
Again, we assume that $r_{1}$ is a single root and $r_{2}$ is a double root of the cubic equation. Then $r_{1}$ and $A_{\mu}$, $\mu=1,2$, are given by (6.77) and (6.78), respectively. Now the linear system is solved by
@@
$C=-\frac{1}{\alpha_{1}\alpha_{2}}\,\left(\begin{array}[]{cccc}0&0&0&0\\ \tilde{r}\,\gamma&0&0&0\\ \frac{1}{4}\big{(}B_{1}v_{4}+2(v_{4}v_{6}+v_{5})\,\tilde{r}\gamma\big{)}v_{4}^{-2}\gamma^{-2}&\frac{1}{4}v_{4}^{-1}v_{5}\tilde{r}\gamma^{-2}&0&0\\ \frac{1}{2}B_{2}v_{4}^{-2}\gamma^{-1}&\frac{1}{4}B_{1}v_{4}^{-1}\gamma^{-2}&\tilde{r}\,\gamma&0\end{array}\right)\,$,
@@
$\tilde{r}$ $=$ $\mathrm{i}\,(\alpha_{1}+\alpha_{2})-2r_{2}\,$,
@@
$f_{2}$ $=$ $-\frac{c_{2}\tilde{r}\gamma}{\alpha_{1}\alpha_{2}}\,t+c_{3}\,$,
$f_{3}$ $=$ $\frac{1}{8(\alpha_{1}\alpha_{2}v_{4}\gamma)^{2}}\Big{(}c_{2}v_{4}v_{5}\tilde{r}^{2}\gamma\,t^{2}-2\alpha_{1}\alpha_{2}\big{(}((2c_{2}v_{6}\gamma+c_{3}v_{5})\,v_{4}+2c_{2}v_{5}\gamma)\,\tilde{r}+c_{2}B_{1}v_{4}\big{)}\,t$
@@
$f_{4}$ $=$ $\frac{1}{24(\alpha_{1}\alpha_{2})^{3}(v_{4}\gamma)^{2}}\Big{(}-c_{2}v_{4}v_{5}\tilde{r}^{3}\gamma^{2}\,t^{3}+3\big{(}(2(v_{4}v_{6}+v_{5})\,c_{2}\gamma+c_{3}v_{4}v_{5})\,\tilde{r}+2c_{2}B_{1}v_{4}\big{)}\,\alpha_{1}\alpha_{2}\tilde{r}\gamma\,t^{2}$
$-6\alpha_{1}^{2}\alpha_{2}^{2}\,(2c_{2}B_{2}\gamma+c_{3}B_{1}v_{4}+4c_{4}v_{4}^{2}\tilde{r}\gamma^{3})\,t+24c_{5}(\alpha_{1}\alpha_{2})^{3}(v_{4}\gamma)^{2}\Big{)}\,$,
@@ 7 Conclusions
We derived a vectorial binary Darboux transformation (bDT) for matrix versions of the first negative flow of the Kaup-Newell (KN) hierarchy. For a reduction of the latter system to matrix generalizations (see (1.2)) of the Fokas-Lenells (FL) equation, we found a corresponding reduction of the bDT to a vectorial Darboux transformation of the latter and generated matrix soliton solutions of (1.2) from the trivial seed solution $u=0$.
Furthermore, we explored in detail the vectorial Darboux transformation with a plane wave seed solution in the case of the two-component vector FL equation, recovering in a different and more systematic way classes of solutions obtained previously in [7, 10, 13]. But we also presented solutions that have not been obtained before to the best of our knowledge. In our approach, black solitons emerge in cases where the matrix $\Gamma$ in the Lyapunov equation does not satisfy the spectrum condition, so that the latter does not have a unique solution. Rogue waves appear in cases where the constant coefficient linear system is “degenerate”, then having quasi-rational solutions.
An interesting feature, known to show up in vector NLS equations, is “beating solitons” (see [26, 27, 28, 23], and references cited there). We presented corresponding solutions for the two-component vector FL equation in Examples 6.2 and 6.3, apparently for the first time. As in the cited work on beating soliton solutions of vector NLS equations, here the relative wave number has been restricted to zero. Similar results have been obtained in [26] for the Manakov system (i.e., the two-component vector NLS equation), also see [28, 23].
@@
Appendix A: Special examples from the class of solutions in Example 6.2
@@
$u_{2}^{\prime}$ $=$ $-\frac{2\,e^{t+x}\,(e^{(1+\mathrm{i}\,)t+\mathrm{i}\,x}+\mathrm{i}\,e^{x})}{(1+\mathrm{i}\,)\,e^{2t}+(1+\mathrm{i}\,)\,e^{2x}+2\,e^{(1+\mathrm{i}\,)(t+x)}+e^{2(t+x)}}\,.$ (A.1)
@@
$u_{1}^{\prime}=e^{\mathrm{i}\,(t+x)}\,\Big{(}1-\frac{2\,e^{-x}}{e^{x}+(1+\mathrm{i}\,)\,e^{-x}}\Big{)}\,,\qquad u_{2}^{\prime}=-\frac{2\,e^{\mathrm{i}\,(t+x)}}{e^{x}+(1+\mathrm{i}\,)\,e^{-x}}\,.$ (A.2)

### v2

@@ 6.2.4 The double root case
Let $n=1$. If the cubic characteristic equation (6.62) admits a single root $r_{1}$ and a double root $r_{2}$, then it is equivalent to
@@
The imaginary parts vanish if and only if $\mathrm{Re}(r_{2})=0$. Hence
$r_{2}=\mathrm{i}\,s\,,\qquad s\in\mathbb{R}\,$,
@@
$|A_{1}|^{2}$ $=$ $\frac{(\alpha_{2}-2s)\,(2ks-1-2\alpha_{1}k)^{2}}{4\alpha_{1}^{2}\,k\,(\alpha_{2}-\alpha_{1})}\,,\qquad|A_{2}|^{2}=-\frac{(\alpha_{1}-2s)\,(2ks-1-2\alpha_{2}k)^{2}}{4\alpha_{2}^{2}\,k\,(\alpha_{2}-\alpha_{1})}\,$,
@@
$\frac{\alpha_{2}-2s}{k\,(\alpha_{2}-\alpha_{1})}>0\,,\qquad\frac{\alpha_{1}-2s}{k\,(\alpha_{2}-\alpha_{1})}<0\,$,
and implies $(\alpha_{1}-2s)(\alpha_{2}-2s)<0$. The Lyapunov equation can only be satisfied if either $c_{1}=0$ or $c_{2}=c_{3}=0$.
@@
$(A^{\dagger}\boldsymbol{\alpha}A)^{2}\,\Big(|c_{3}|^{2}+\frac{\|A\|^{4}\,|\tilde{c}|^{2}}{|A_{1}|^{2}\,|A_{2}|^{2}\,(\alpha_{2}-\alpha_{1})^{2}}\Big)=k\,\|A\|^{2}\,\Big|\mathrm{i}\,c_{2}-c_{3}\,s+c_{3}\,\big(\frac{1}{2k}+\frac{A^{\dagger}\boldsymbol{\alpha}A}{\|A\|^{2}}\big)-\tilde{c}\,\Big|^{2}\,$,
@@
$\tilde{c}$ $=$ $\Big((1-A^{\dagger}\boldsymbol{\alpha}A)\,k^{-1}+\frac{\alpha_{1}\alpha_{2}\|A\|^{2}}{A^{\dagger}\boldsymbol{\alpha}A}\Big)^{-1}\Big(2\mathrm{i}\,c_{2}\,s-c_{3}\,s^{2}-\mathrm{i}\,\frac{A^{\dagger}\boldsymbol{\alpha}^{2}A}{A^{\dagger}\boldsymbol{\alpha}A}(c_{2}+\mathrm{i}\,c_{3}\,s)+\frac{1}{4}k^{-2}c_{3}$
@@
$\Omega$ $=$ $\frac{\|A\|^{2}}{k(A^{\dagger}\boldsymbol{\alpha}A)^{2}\,\varrho_{2}^{2}}\Big(\frac{\mathrm{i}\,k\,\tilde{s}\,|c_{2}|^{2}}{3\alpha_{1}\alpha_{2}}\varrho_{3}\,(\mathrm{i}\,A^{\dagger}\boldsymbol{\alpha}A\,\varrho_{2}-\varrho_{3})\,(3xt^{2}-\frac{3\mathrm{i}\,\alpha_{1}\alpha_{2}}{k\tilde{s}}x^{2}t+\frac{\mathrm{i}\,k\tilde{s}}{\alpha_{1}\alpha_{2}}t^{3})$
$-\mathrm{i}\,\big(k|c_{2}|^{2}A^{\dagger}\boldsymbol{\alpha}A\,\varrho_{2}(\varrho_{2}-k\varrho_{1})+2\mathrm{i}\,\mathrm{Re}(c_{2}c_{3}^{\ast})\,\varrho_{3}(\mathrm{i}\,A^{\dagger}\boldsymbol{\alpha}A\,\varrho_{2}-\varrho_{3})\big)\,(xt+\frac{\mathrm{i}\,k\tilde{s}}{2\alpha_{1}\alpha_{2}}t^{2})$
@@
$+\mathrm{i}\,k|c_{2}|^{2}A^{\dagger}\boldsymbol{\alpha}A\,\|A\|^{2}\varrho_{4}^{2}\,(k\varrho_{1}-\varrho_{2})\,\big(\mathrm{i}\,ks\,(k\varrho_{1}-\varrho_{2})+\varrho_{3}\big)+\varrho_{8}\big)\,x+c\Big)\,$,
@@
$\tilde{s}$ $=$ $\mathrm{i}\,(\alpha_{1}+\alpha_{2}-2s)\,$,
$\varrho_{1}$ $=$ $-2\,s\,A^{\dagger}\boldsymbol{\alpha}A+A^{\dagger}\boldsymbol{\alpha}^{2}A\,$,
@@
$\varrho_{3}$ $=$ $-\mathrm{i}\,k^{2}\,\big(A^{\dagger}\boldsymbol{\alpha}A\,s^{2}-(\alpha_{1}\alpha_{2}\,\|A\|^{2}+A^{\dagger}\boldsymbol{\alpha}^{2}A)\,s+\alpha_{1}\alpha_{2}\,A^{\dagger}\boldsymbol{\alpha}A\big)$
$-\mathrm{i}\,k\,\big(s\,(A^{\dagger}\boldsymbol{\alpha}A)^{2}+\frac{1}{2}\alpha_{1}\alpha_{2}\|A\|^{2}+\frac{1}{2}\varrho_{1}\big)+\frac{1}{4}\mathrm{i}\,A^{\dagger}\boldsymbol{\alpha}A\,(2A^{\dagger}\boldsymbol{\alpha}A-1)\,$,
$\varrho_{4}$ $=$ $\frac{1}{2k^{2}}\Big(\|A\|^{2}\,\big(2\mathrm{i}\,(\mathrm{i}\,ks\,\varrho_{2}+\varrho_{3})+\varrho_{2}\big)+2kA^{\dagger}\boldsymbol{\alpha}A\,\varrho_{2}\Big)\,$,
$\varrho_{5}$ $=$ $(A^{\dagger}\boldsymbol{\alpha}A)^{2}\varrho_{2}\,(1-2ks)\,(k\varrho_{4}-A^{\dagger}\boldsymbol{\alpha}A\,\varrho_{2})-2\alpha_{1}\alpha_{2}\|A\|^{2}(k(A^{\dagger}\boldsymbol{\alpha}A)^{2}\varrho_{2}^{2}+\|A\|^{2}\varrho_{3}^{2})\,$,
$\varrho_{6}$ $=$ $\|A\|^{2}\varrho_{1}\,(1-2ks)-2(k\varrho_{4}-A^{\dagger}\boldsymbol{\alpha}A\,\varrho_{2})\,$,
$\varrho_{7}$ $=$ $(\alpha_{1}\alpha_{2}\|A\|^{2}-s\,A^{\dagger}\boldsymbol{\alpha}A)\,(k\,(A^{\dagger}\boldsymbol{\alpha}A)^{2}\varrho_{2}^{2}+\|A\|^{2}\varrho_{3}^{2})\,$,
$\varrho_{8}$ $=$ $|c_{2}|^{2}A^{\dagger}\boldsymbol{\alpha}A\,\|A\|^{4}\varrho_{1}\varrho_{3}^{2}\varrho_{4}+\frac{1}{2}\varrho_{4}\,\Big(4\mathrm{i}\,\mathrm{Im}(c_{2}c_{3}^{\ast})\,\big(k\,s\,A^{\dagger}\boldsymbol{\alpha}A\,\|A\|^{2}\varrho_{3}\varrho_{4}\,(k\varrho_{1}-\varrho_{2})+\mathrm{i}\,\|A\|^{2}\varrho_{1}\varrho_{7}\big)$
$+\mathrm{i}\,k\,c_{2}^{\ast}c_{3}\,(A^{\dagger}\boldsymbol{\alpha}A)^{2}\|A\|^{2}\varrho_{1}\varrho_{2}\varrho_{4}\,(1-2ks)-|c_{3}|^{2}\varrho_{4}\varrho_{5}\Big)\,$,
@@
Figure 14: Plots of the absolute values of the two components of a “double pole solution” from the class in Example 6.18 (2). Here we set $\alpha_{1}=1,\alpha_{2}=2,k=1,s=3/4,c_{1}=0,c_{2}=1+\mathrm{i}$ and $c_{3}=1+(353/1365)\,\mathrm{i}$, so that the constraint from the Lyapunov equation is satisfied.
@@
Next we present a second-order version of the above class of solutions in Example 6.18.
@@
Again, we assume that $r_{1}$ is a single root and $r_{2}$ is a double root of the cubic equation (6.62). Then $r_{1}$ and $A_{\mu}$, $\mu=1,2$, are given by (6.93) and (6.94), respectively. Now the linear system is solved by
@@
$C=-\frac{1}{\alpha_{1}\alpha_{2}}\,\left(\begin{array}[]{cccc}0&0&0&0\\ \tilde{s}\,\gamma&0&0&0\\ \frac{1}{4}\big(B_{1}v_{4}+2(v_{4}v_{6}+v_{5})\,\tilde{s}\gamma\big)v_{4}^{-2}\gamma^{-2}&\frac{1}{4}v_{4}^{-1}v_{5}\tilde{s}\gamma^{-2}&0&0\\ \frac{1}{2}B_{2}v_{4}^{-2}\gamma^{-1}&\frac{1}{4}B_{1}v_{4}^{-1}\gamma^{-2}&\tilde{s}\,\gamma&0\end{array}\right)\,$,
@@
$\tilde{s}$ $=$ $\mathrm{i}\,(\alpha_{1}+\alpha_{2})-2r_{2}\,$,
@@
$f_{2}$ $=$ $-\frac{c_{2}\tilde{s}\gamma}{\alpha_{1}\alpha_{2}}\,t+c_{3}\,$,
$f_{3}$ $=$ $\frac{1}{8(\alpha_{1}\alpha_{2}v_{4}\gamma)^{2}}\Big(c_{2}v_{4}v_{5}\tilde{s}^{2}\gamma\,t^{2}-2\alpha_{1}\alpha_{2}\big(((2c_{2}v_{6}\gamma+c_{3}v_{5})\,v_{4}+2c_{2}v_{5}\gamma)\,\tilde{s}+c_{2}B_{1}v_{4}\big)\,t$
@@
$f_{4}$ $=$ $\frac{1}{24(\alpha_{1}\alpha_{2})^{3}(v_{4}\gamma)^{2}}\Big(-c_{2}v_{4}v_{5}\tilde{s}^{3}\gamma^{2}\,t^{3}+3\big((2(v_{4}v_{6}+v_{5})\,c_{2}\gamma+c_{3}v_{4}v_{5})\,\tilde{s}+2c_{2}B_{1}v_{4}\big)\,\alpha_{1}\alpha_{2}\tilde{s}\gamma\,t^{2}$
$-6\alpha_{1}^{2}\alpha_{2}^{2}\,(2c_{2}B_{2}\gamma+c_{3}B_{1}v_{4}+4c_{4}v_{4}^{2}\tilde{s}\gamma^{3})\,t+24c_{5}(\alpha_{1}\alpha_{2})^{3}(v_{4}\gamma)^{2}\Big)\,$,
@@ 7 Conclusions
We derived a vectorial binary Darboux transformation (bDT) for matrix versions of the first negative flow of the Kaup-Newell (KN) hierarchy. For a reduction of the latter system to matrix generalizations (see (1.2)) of the Fokas-Lenells (FL) equation, we found a corresponding reduction of the bDT to a vectorial Darboux transformation of the latter and generated various classes of matrix soliton solutions of (1.2) from the trivial seed solution $u=0$.
Furthermore, we explored in detail the vectorial Darboux transformation with a plane wave seed solution in the case of the two-component vector FL equation, recovering in a different and more systematic way classes of solutions obtained previously in [7, 11, 16]. But we also presented solutions that have not been obtained before to the best of our knowledge. This concerns in particular solutions corresponding to cases, where the matrix $\Gamma$ in the Lyapunov equation does not satisfy the spectrum condition, so that the Lyapunov equation does not have a unique solution.
An interesting feature, known to show up in vector NLS equations, is “beating solitons” (see [34, 35, 36, 32], and references cited there). We presented corresponding solutions for the two-component vector FL equation in Examples 6.5 and 6.6, apparently for the first time. As in the cited work on beating soliton solutions of vector NLS equations, here the relative wave number has been restricted to zero. Similar results have been obtained in [34] for the Manakov system (i.e., the two-component vector NLS equation), also see [36, 32].
@@
We would like to highlight again the nonlinear superposition property of a vectorial Darboux transformation, for solutions generated from the same seed solution, as formulated in Remarks 4.4 and 5.6 for the specific integrable equations studied in this work (also see Appendix Appendix D: Superposition of two solutions from the class in Example 6.5 for a simple example). In our opinion, this is a major advantage over the classical iterative Darboux transformation method.
Finally, we have found further classes of integrable equations that fit into the bidifferential calculus framework.
@@
Appendix A: Some facts about the Lyapunov equation
We consider the Lyapunov equation
$\Gamma\,\Omega+\Omega\,\Gamma^{\dagger}=K\,$, (A.1)
with $n\times n$ matrices $\Gamma,\Omega,K$. Here we think of $\Gamma$ and $K$ as given matrices and look for a solution $\Omega$.
Let $\Gamma=\mathrm{diag}(\gamma_{1},\ldots,\gamma_{n})$, with $\gamma_{i}\neq-\gamma_{j}^{\ast}$, $i,j=1,\ldots,n$. Then (A.1) reads
$(\gamma_{i}-\gamma_{j}^{\ast})\,\Omega_{ij}=K_{ij}\qquad i,j=1,\ldots,n\,.$
If $\Gamma$ satisfies the spectrum condition, the Lyapunov equation possesses a unique solution, given by the Cauchy-like matrix
$\Omega=\left(\frac{K_{ij}}{\gamma_{i}+\gamma_{j}^{\ast}}\right)\,.$
If, for some values of $i$ and $j$, $\gamma_{i}=\gamma_{j}^{\ast}$, then a solution can only exist if the corresponding component $K_{ij}$ of $K$ vanishes, and $\Omega_{ij}$ is left undetermined.
Let $\Gamma$ be the lower triangular $n\times n$ Jordan block
$\Gamma=\left(\begin{array}[]{ccccc}\gamma&0&\cdots&\cdots&0\\ 1&\gamma&\ddots&\ddots&0\\ 0&\ddots&\ddots&\ddots&\vdots\\ \vdots&\ddots&\ddots&\ddots&0\\ 0&\cdots&0&1&\gamma\end{array}\right)\,$, (A.7)
with $\gamma\in\mathbb{C}$. If the spectrum condition holds, we have the following result [33].
Proposition A.1.
Let $\mathrm{Re}(\gamma)\neq 0$. For $1\leq k<n$, let $K_{(k)}$ and $\Gamma_{(k)}$ be the $k\times k$ upper left part of $K$, respectively $\Gamma$, and $\Omega_{(k)}$ the solution of
$\Gamma_{(k)}\,\Omega_{(k)}+\Omega_{(k)}\,\Gamma_{(k)}^{\dagger}=K_{(k)}\,.$
For $k=1,\ldots,n-1$, we have
$\Omega_{(k+1)}=\left(\begin{array}[]{cc}\Omega_{(k)}&B_{(k)}\\ B_{(k)}^{\dagger}&\omega_{(k+1)}\end{array}\right)\,$,
where
$B_{(k)}:=\hat{\Gamma}_{(k)}^{-1}\,\Big((K_{1,k+1},K_{2,k+1},\ldots,K_{k,k+1})^{T}-\Omega_{(k)}(0,\ldots,0,1)^{T}\Big)\,$,
$\omega_{(k+1)}:=\frac{1}{\kappa}\Big(K_{k+1,k+1}-2\,\mathrm{Re}[(0,\ldots,0,1)B_{(k)}]\Big)\,$,
and $\hat{\Gamma}_{(k)}$ is the Jordan block $\Gamma_{(k)}$ with $\gamma$ replaced by $\kappa=2\,\mathrm{Re}(\gamma)$. $\Box$
This proposition allows to recursively compute the solution of the Lyapunov equation (A.1) with a Jordan matrix $\Gamma_{(n)}$. The inverse of $\Omega_{(n)}$ can also be recursively computed via
$\Omega_{(k+1)}^{-1}=\left(\begin{array}[]{cc}\Omega_{(k)}^{-1}+S_{\Omega_{(k)}}^{-1}\Omega_{(k)}^{-1}B_{(k)}B_{(k)}^{\dagger}\Omega_{(k)}^{-1}&-S_{\Omega_{(k)}}^{-1}\Omega_{(k)}^{-1}B_{(k)}\\ -S_{\Omega_{(k)}}^{-1}B_{(k)}^{\dagger}\Omega_{(k)}^{-1}&S_{\Omega_{(k)}}^{-1}\end{array}\right)\,$, (A.11)
with the scalar Schur complement (a quasi-determinant)
$S_{\Omega_{(k)}}=\omega_{(k+1)}-B_{(k)}^{\dagger}\Omega_{(k)}^{-1}B_{(k)}\,.$
Example A.2.
For $n=2$, we obtain
$\Omega_{(2)}=\frac{1}{\kappa}\left(\begin{array}[]{cc}K_{11}&K_{12}-\kappa^{-1}K_{11}\\ K_{21}-\kappa^{-1}K_{11}&K_{22}-\kappa^{-1}(K_{12}+K_{21})+2\kappa^{-2}K_{11}\end{array}\right)\,.$
Its determinant is
$\det(\Omega_{(2)})=\kappa^{-4}\,K_{11}^{2}+\kappa^{-2}\,\det(K)\,.$ (A.13)
$\Box$
Proposition A.3.
Let $\Gamma$ be the $n\times n$ lower triangular Jordan block with imaginary eigenvalue $\gamma$. Then the Lyapunov equation (A.1) is equivalent to
$K_{11}=0\,$,
$\Omega_{i-1,j}=K_{i1}\qquad i=2,3,\ldots,n\,$,
$\Omega_{1,j-1}=K_{1j}\qquad i=2,3,\ldots,n\,$,
$\Omega_{i-1,j}+\Omega_{i,j-1}=K_{ij}\qquad i,j=2,3,\ldots,n\,.$
We note that $\Omega_{nn}$ drops out of the Lyapunov equation.
Proof.
Expressed in components, the Lyapunov equation reads
$\sum_{k=1}^{n}\Big(\Gamma_{ik}\Omega_{kj}+\Omega_{ik}\Gamma^{\dagger}_{kj}\Big)=\Gamma_{ii}\Omega_{ij}+\Gamma_{i,i-1}\Omega_{i-1,j}+\Omega_{ij}\Gamma^{\dagger}_{jj}+\Omega_{i,j-1}\Gamma^{\dagger}_{j-1,j}=K_{ij}\,$,
where $i,j=1,\ldots,n$. This is equivalent to
$(\gamma+\gamma^{\ast})\,\Omega_{11}=K_{11}\,$,
$(\gamma+\gamma^{\ast})\,\Omega_{i1}+\Omega_{i-1,j}=K_{i1}\qquad i=2,3,\ldots,n\,$,
$(\gamma+\gamma^{\ast})\,\Omega_{1j}+\Omega_{1,j-1}=K_{1j}\qquad i=2,3,\ldots,n\,$,
$(\gamma+\gamma^{\ast})\,\Omega_{ij}+\Omega_{i-1,j}+\Omega_{i,j-1}=K_{ij}\qquad i,j=2,3,\ldots,n\,.$
If $\gamma$ is imaginary, this reduces to the equations in the proposition. ∎
Example A.4.
For the $2\times 2$ Jordan block with imaginary eigenvalue, the Lyapunov equation imposes the constraints $K_{11}=0$, $K_{21}=K_{12}$, and only determines part of $\Omega$ via $\Omega_{11}=K_{12}$ and $\Omega_{12}+\Omega_{21}=K_{22}$. $\Box$
Appendix B: Special examples from the class of solutions in Example 6.5
@@
$u_{2}^{\prime}$ $=$ $-\frac{2\,e^{t+x}\,(e^{(1+\mathrm{i})\,t+\mathrm{i}\,x}+\mathrm{i}\,e^{x})}{(1+\mathrm{i})\,e^{2t}+(1+\mathrm{i})\,e^{2x}+2\,e^{(1+\mathrm{i})(t+x)}+e^{2(t+x)}}\,.$ (B.1)
@@
$u_{1}^{\prime}=e^{\mathrm{i}\,(t+x)}\,\Big(1-\frac{2\,e^{-x}}{e^{x}+(1+\mathrm{i})\,e^{-x}}\Big)\,,\qquad u_{2}^{\prime}=-\frac{2\,e^{\mathrm{i}\,(t+x)}}{e^{x}+(1+\mathrm{i})\,e^{-x}}\,.$ (B.2)